Circle the number of the plot showing the direction field for each of the following differential equations.

(1) y′ = x + y − 1 I II III None (2) y′ = 2 − y I II III None (3) y′ = x(2 − y) I II III None (4) y′ = sin(x) sin(y) I II III None

I: II:

III: